The fish swims 2 m below the flat surface of the lake. At what apparent depth below the surface is the fish for the observer directly above it on the surface of the lake? The refractive index of water is 1.33.


1
Expert's answer
2021-04-22T11:02:22-0400

"\mu_{water} = \frac{real \; depth}{apparent \; depth}"

apparent depth "= \frac{real \; depth}{\mu_{water}}"

"= \frac{2.00}{1.33} = 1.50 \;m"

Answer: 1.50 m